Fermi National Laboratory


Accelerator Update for October 12, 2005 - October 14, 2005

Wednesday October 12
The day shift began with Operations monitoring store 4442 and a Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with NuMI off while its Horn ground fault repairs finish, and with MiniBooNE and the Meson MCenter experiments taking beam.

Operations began preparing the Main Injector (MI) at 10:19 AM for an access to work on a NuMI kicker. The stash was dumped for this access.

At 11:01 AM, Operations reported a vacuum cascade in Tevatron (TeV) sector F0. Operators quickly reset vacuum pumps, but the luminosity still took a hit.

Technicians reported at 12:07 PM that they had completed their work on the NuMI kicker.

Operations reported at 12:31 PM that the MI access had ended.

Operations established beam to MiniBooNE at 1:15 PM.

Pbar began stacking at 1:37 PM.

MI experts began a study period at 2:21 PM. Unfortunately, it was causing problems in Booster due to some Time Line Generator (TLG) conflicts. Experts investigated.

At 2:22 PM, a TeV air compressor (T:D1AIRP) tripped off. Operators responded and reset the compressor. While there, they noticed a small pressure leak. They notified experts who connected a portable compressor to the system. The experts will wait to repair the compressor until after the store gets terminated.

CDF experimenters reported at 3:26 PM that they were starting to power up their solenoid. The solenoid had tripped off a few days ago due to a ruptured helium relief valve.

MI studies ended at 4:28 PM. Operations began preparing the Recycler to stash antiprotons. The stash ended around 6:10 PM.

MCenter resumed taking beam at 6:29 PM.

At 7:06 PM, Operations reported a problem with the NuMI target cooling at MI-65, which was holding off the beam permit. Operations contacted an expert and had the cooling problem resolved by 7:32 PM.

Operations reported at 8:28 PM that NuMI had received its first pulse of beam since the Horn repairs.

Pbar resumed stacking at 10:50 PM.

MCenter resumed taking beam at 11:02 PM.

Thursday October 13
The midnight shift began with Operations monitoring store 4442 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:35 AM, a Booster RF station (BRF15) tripped off for the first of two times in two minutes. It reset okay.

At 12:59 AM, BRF7 & 8 tripped off on a vacuum pressure indication. The stations wouldn't reset. Operators found a vacuum pump off. They reset the pump and then reset the stations.

At 1:40 AM, a NuMI power supply (E:QD116) tripped off on a phase loss indication. This trip took down the NuMI beam permit. The supply and permit reset okay.

At 2:14 AM, the E:QD116 supply and NuMI beam permit tripped off again. The supply wouldn't reset. Operators changed its control card, but this didn't immediately resolve the problem. Operations contacted a power supply expert. Operators got the supply on at 2:45 AM, but it tripped off again five minutes later. This tripping continued for a few minutes and then Operations called in an expert. The power supply technician cleaned the breaker at 3:53 AM, but this didn't solve the problem. He changed the supply's contactor at 4:47 AM.

At 4:54 AM, E:QD116 tripped off. This time it took down the MI PSLOOP. Operations reset it, but it tripped off again a few minutes later. Operations put in a stacking only timeline.

TeV experts began an end-of-store study around 6:30 AM.

At 6:50 AM, Operations tried a NuMI/stacking timeline, but QD116 tripped. Operations returned to a stacking only timeline.

The power supply expert was ready to try operating QD116 at 8:24 AM. He had replaced the SCR bridge and this fixed the problem. NuMI resumed taking beam at 8:25 AM.

Operations terminated store 4442 at 9:15 AM.

Cryo system techs began cryo maintenance work on TeV sectors B1 & 2 at 9:30 AM.

Operations reestablished beam to MCenter at 9:49 AM.

Operations prepared Booster for an access at 11:38 AM. A Booster expert inspected a kicker (MKS06).

D0 came out of access at 12:03 PM.

Operations reported at 1:16 PM that the Pelletron's Electron Cooling repairs had been completed and the Pelletron had resumed cooling the stash.

At 1:28 PM, a Booster expert reported that MKS06 was failing. Experts repaired the kicker magnet and were out of the tunnel by 2:30 PM.

Pbar resumed stacking at 3:03 PM.

Operations put the TeV through a dry squeeze at 3:09 PM.

MCenter resumed taking beam at 3:34 PM.

Operations put the TeV through a wet squeeze at 4:09 PM.

Operations put the TeV into shot setup at 4:40 PM.

Operations established combination store 4445 at 7:06 PM with an initial luminosity of 142.51E30. This store set a NEW LUMINOSITY RECORD. Pbar resumed stacking. MCenter resumed taking beam.

Recycler began preparing to stash at 9:14 PM. The stash ended and Pbar resumed stacking at 10:03 PM. MCenter resumed taking beam at 10:13 PM.

At 10:32 PM, a TeV kicker power supply (C:PAK5) faulted. Operators investigated and contacted an expert. Using the expert's advice, operators had the kicker supply up within five minutes.

Friday October 14
The midnight shift began with Operations monitoring store 4445 and the stash, with Pbar stacking, and with NuMI, MiniBooNE and MCenter taking beam.

Recycler began preparing to stash at 2:30 AM. The stash ended at 3:13 AM. Pbar and MCenter resumed normal operations.

At 6:33 AM, the Tevatron Electron Lens (TEL) tripped off. It reset okay.

Recycler began preparing to stash at 7:24 AM. The stash ended at 8:07 AM. Pbar and MCenter resumed normal operations.

The Plans for Friday and the Weekend
The plans for today are to stack, stash, and store. The next store will be a Recycler only store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/14/2005   email Fermilab